OpenStudy (anonymous):

Find a solution to: y' = 1 + ytan x when y(0) = 1

OpenStudy (anonymous):

?

myininaya (myininaya):

\[y'-\tan(x) y=1\]

myininaya (myininaya):

we want to multiply both sides by \[v(x)=e^{\int\limits_{}^{}tanx dx}=e^{\int\limits_{}^{}\frac{\sin(x)}{\cos(x)} dx} \] let \[u=\cos(x)=> du=-\sin(x) dx\] \[v(x)=e^{\int\limits_{}^{}\frac{-1}{u} du}=e^{-\ln(u)}=e^{-\ln(\cos(x))}=\frac{1}{\cos(x)}\]

myininaya (myininaya):

\[\frac{1}{\cos(x)}y'-\frac{1}{\cos(x)}\tan(x)y=\frac{1}{\cos(x)}\] \[\sec(x)y'-\sec(x)\tan(x)y=\sec(x)\] we can write \[(\sec(x) y)'=\sec(x)\]

myininaya (myininaya):

now integrate both sides

myininaya (myininaya):

\[\sec(x) y=\int\limits_{}^{}\sec(x) \cdot \frac{\sec(x)+\tan(x)}{\sec(x)+\tan(x)} dx +C\] \[\sec(x) y=\int\limits_{}^{}\frac{du}{u}+C=\sec(x) y=\ln|u|+C=\ln|\sec(x)+\tan(x)|+C\]

myininaya (myininaya):

so we have \[\sec(x) y=\ln|\sec(x)+\tan(x)|+C\] now when x=0,y=1 \[\sec(0)(1)=\ln|\sec(0)+\tan(0)|+C\]

myininaya (myininaya):

solve for C

myininaya (myininaya):

\[(1)(1)=\ln|(1)+(0)|+C\]

myininaya (myininaya):

\[1=0+C\]

myininaya (myininaya):

\[C=1 => \sec(x) y=\ln|\sec(x)+\tan(x)|+1\]

OpenStudy (anonymous):

\[y =(\ln|\sec(x)+\tan(x)|+1)/\sec(x)\]

OpenStudy (anonymous):

Thanks

myininaya (myininaya):

np

OpenStudy (zarkon):

hmmm...

OpenStudy (zarkon):

I believe the integrating factor should be \[\cos(x)\]

myininaya (myininaya):

i disagree

OpenStudy (zarkon):

\[y'-\tan(x) y=1\] \[y'+(-\tan(x))y=1\] \[u=e^{\int-tan(x)dx}\] \[=\cos(x)\]

OpenStudy (zarkon):

must be in standard form

myininaya (myininaya):

oops i forgot about the negative

myininaya (myininaya):

fine you are right

myininaya (myininaya):

lol

OpenStudy (zarkon):

:)

OpenStudy (zarkon):

I try to be when I can

OpenStudy (zarkon):

makes the problem easier ... that's a good thing

myininaya (myininaya):

\[\cos(x) y'-\cos(x)\tan(x)y=\cos(x)\] \[(\cos(x) y)'=\cos(x)\]

myininaya (myininaya):

\[\cos(x) y=\sin(x)+C\]

myininaya (myininaya):

tons easier!

OpenStudy (zarkon):

yep

myininaya (myininaya):

\[\cos(0)(1)=\sin(0)+C =>C=1\]

myininaya (myininaya):

\[\cos(x) y=\sin(x)+1\]

OpenStudy (anonymous):

Wow, I though those secant integrals seemed a bit unexpected! Relieved